Two bodies of masses m1 and m2 are connected to the ends of a massless cord and allowed to move.pulley-massless and frictionless.find accelaration of centre of mass.

Grade:11

1 Answers

vikas askiitian expert
13 years ago

T - m1g = m1a       ...........1

m2g - T = m2a         ...........2             (m2 > m1 )

adding both

a = (m2-m1)g/(m1+m2)     .............3

 

now , m1 will move upward but m2 will move downward ...

accleration of m1 is a1 = a(j)                    (upward is taken as +ve)

accleration of m2 is a2 = a(-j)                    (downward is taken as -ve)

 

accleration of center of mass is acom = m1a1 + m2a2 /m1+m2

 acom = (m1a - m2a)/m1+m2  (j)

substituting a from eq 3

    acom  = -g(m2-m1)2/(m1+m2)2  (j)

accleration of com will be in downward direction